Summary of the 3mm GMVA Session of May 22-27, 2014 -------------------------------------------------- General: On has a new 3mm receiver, now supporting dual polarisation. Yb uses a borrowed 3mm receiver from IRAM, also with dual polarisation. Eb is recording using the RDBE. Noto is participating at 7mm (jointly with the VLBA). Yb was recording for a small time fraction at 7mm (and fringes EB_RDBE-Yebes_DBBC are found) May 22: fringe test with Eb, On, Yb, PV, PdB part1 - f141a: Fringes on Eb-On, no fringes with Yebes part2 - f141b: PdB and PV not available due to bad weather redo fringe check Eb-Yb -> still no fringes found Pico cannot participate due to a DBBC problem part3 - c141b: Fringes on Eb-PdB in zoom mode (32/16 MHz) General stations comments: -------------------------- Effelsberg - observed the whole session, no major problem except very variable and partly bad weather (rain, thick clouds). Pico Veleta - due to a broken chip in the DBBC they cannot participate in the fringe test f141b. Replacement card sent from Bonn to Granada via UPS, arrives on May 23. Pico joins c141a on May 23, at 17h45. That means they lost the first 24 hrs of c141a (CygA). From then on, Pico observed until the end of the session, however interrupted by bad weather. During c141b the weather was quite good. Note added after fringesearch: only one polarisation, other being overwritten, dbbctrk2 command does not work, DBBC2 Bug in firmware confirmed by G. Tuccari on July 17. It may be related to the known problem setting ADB1/ADB2. PdBure - due to bad weather, late arrival on the mountain (helicopter transport delayed until May 23). First scan recorded May 23, 18h15. So they also lost the first 20 hrs (CygA). From then on observing until end, with several interruption due to bad weather (snow, wind, fog). Reference antenna is No 3 on position N07. Noto - observing, no detailed comments sent Yebes- observing, see below VLBA - NL not participating (log says: scheduled maintenance visit) PT, MK frequently taken out from the GMVA run for Geodetic Observations (several 1 hr blocks per day) PT not observing on May 25, day 145 from 0921-1521 UT , Azimut drive problem Detailed station comments: --------------------------- Effelsberg: day 142 - May 22: started f141a in time, thick clouds in Effelsberg is was raining during all of f141b, Tsys is very high day 143 - May 23: sky is clearing up at 18 UT day 144 - May 24: after a clear night, a rain front passed through, and now (7 UT) the clouds are getting thinner. Tsys=180K, we can see and point on BLLAC. All systems running. day 145 - May 25: For Effelsberg observing is going well, last night was clear, tau=0.09. Now during daytime the cloud coverage has increased, and convection has started. day 146 - May 26: During the 2nd half of the night clouds came in and now it its overcast. From 12UT-18UT we'll switch to 7mm for the 3C84 project. day 147 - May 27: Last night was rainy and not good, and now it is overcast. However, we do see BLLAC and can point on it. Tsys 210K. Onsala: day 142 - May 22: fringetest reveals swapped polarisation at Onsala Yebes: day 142 - May 22: The weather has been very bad until 20:00, Then we have got big clears in the sky. day 143 : The sky was OK, at 2:00 UT, but I have discovered that IFB has had a problem ONLY in C141A (not in f141b and f141a). I have fixed it half an hour ago. So by that time only IFA had good data. day 146 May 26: Yebes observes at 7mm 11h50-18h10 UT (3C84 project, c141b.vex_7mmversion, ver=8) Hope to detect fringes at 3mm and find out if DBBC or receiver has a problem. 19 UT: switch back to 3mm and observe at 3mm until end. Note 16.7.2014: 7mm fringes between EB-Ys are seen, SNR 9156 (3C84) Note 27.8.2014: no fringes at 3mm (possible pointing problem) Pico: day 142 - May 22: broken DBBC board, cannot record data, replacement card will be shipped from Bonn day 143 May 23: could repair the DBBC with the spare part received from Bonn. Have started recording at 17:45 UT. The weather is not very good but sources are well visible. Tsys about 115k. day 144 May 24: The wind was increasing along the night and we stopped one hour ago (6h30UT) Resume observation 11h15 UT (weather still bad, clouds) day 145 May 25: We have changed to part B with very good weather. Tsys about 100k. 23 UT: We have been observing with good conditions for the past 10+ hours, weather forecast is good. day 146 May 26: observing all the night in good weather conditions day 147 May 27: Excellent weather in PV and we are observing without interruption. PdBure: day 143 - May 23: team could reach Plateau de Bure today and start VLBI observations at approx. 10:00 UT under marginal conditions (thick clouds). Had to stop at 11 UT due to high winds. May 23: 22 UT -> correlator confirms fringes for PdB-Ef (1 Gbps) May 24: Plateau de Bure had to stop from 21:55 UT to 03:25 UT last night due to fog, wind and some snow. We observed under excellent weather until about 10:00 UT when convective clouds reached the level of the Plateau. We are observing now under instable conditions. May 25: Bure observed all night and the change to c141b went fine. Currently we have a Tsys of about 90K and stable phases. 20 UT: stop, snowfall and bad weather May 26: resume observation 03h30 UT, We have dense fog (tendency towards improvement) and stable phases. The system is working fine. May 27: after a series of thunderstorms yesterday afternoon, we could only rejoin the session from 21:00 UT on, until 00:36 UT when snow stared to fall. We have restarted observations at 9:00 UT and are observing since in fog. VLBA: VERY LONG BASELINE ARRAY OBSERVING LOG -------------------------------------- Project: C141A Observer: Krichbaum, T. Project type: GEO Obs filename: c141acrd.* Date/Day: 2014MAY23/143 Ants Scheduled: NL FD LA PT KP OV BR MK =UT-Time===Comment===============================================MF#===%AD==AMD= 0156 Begin file VCLOCK OK (OV: -0.023); UT1-UTC OK (-0.28) GPS delta < 10us (MK: -3.5us) Operator: T. Perreault 0156 NL not observing; scheduled maintenance visit SCH (C134706) 0156 %FD wea: light rain 0156-0435 %BR 3mm pointing may not be accurate due to FOCUS FRM 50 1519 motor replacement (C134673) 0206 MK sporadic DBE1-TimestampSyncError alerts are being generated. Short of stopping observation to reboot DBE1 there is little that operations can do. Unknown if this issue is causing problem with recording data onto the pack; there is an open maintenance form on this problem and we're waiting to hear back. Will see if the problem continues when MK is removed for the USNO observation or see if it returns when MK resumes this observation. 0439 KP wea: windy 0455-0550 PT, MK removed for USNO observation SCH 200 110 MK when USNO started the DBE1-TimestampSyncError alert went away as expected since USNO used DBE-2. 0535 Operator is A. Sowinski. 0648-0754 *PT,MK removed for USNO reobs. SCH 200 132 1345 New operator is Alan Kerr 1759-1802 *PT out of point. power glitch. COM 100 3 2145 New operator: T. Perreault 0000 Date is 2014MAY24/144. 0055 %LA wea: raining 0225-0316 *PT, MK removed for USNO observation SCH 200 102 0331 Going to have to re-run the USNO observation; no data was recorded on the USNO packs at either site. Don't know why not since all indications were that it was recording. Set next USNO run to go from 0555 to 0646 UT. 0340-0342 *PT, MK rebooting legacy computer to read in new SCH 200 4 observ.tx file for re-reun of USNO observation. 0538 Operator is A. Sowinski. 0552-0650 *PT,MK removed for USNO observation N4144B. SCH 200 116 1345 New operator is Alan Kerr 1554-1556 *PT out of point. power glitch. COM 100 2 2024-2115 %LA precipitation. 0000 Date is 2014MAY25/145. 0124-0126 *PT autostowed, power glitch, repointed COM 100 2 0130 New operator: T. Perreault 0130 %PT wea: raining 0225-0321 *PT, MK removed for USNO observation N4145 SCH 200 112 0435 End. Downtime Summary: Total downtime : 2102 min Percentage downtime of observing: 8.6% Average downtime per hour : 5.2 min Total scheduled observing time (# Antennas): 24312 min (8) VERY LONG BASELINE ARRAY OBSERVING LOG -------------------------------------- Project: C141B Observer: Krichbaum, T. Project type: GEO Obs filename: c141bcrd.* Date/Day: 2014MAY25/145 Ants Scheduled: FD LA PT KP OV BR MK =UT-Time===Comment===============================================MF#===%AD==AMD= 0441 Begin file VCLOCK OK (LA: -0.041); UT1-UTC OK (-0.28) GPS delta < 10us (MK: -3.5us) Operator: T. Perreault 0441 NL not observing; scheduled maintenance visit SCH (C134706) 0441-2100 %BR 3mm pointing may not be as accurate due to FOCUS FRM motor replacement (C134673) 0825-0919 *PT, MK removed for re-observation of USNO SCH 200 108 0919-1521 &PT unclearable AZ1 armature overcurrent error ACU 100 362 Called oncall site tech, he tried several things from home but was also unable to clear the error condition. If he could have reached the site and gotten the error condition cleared before the window for USNO observations closed for today (1030 UT) I would have had him go out. Instead he will go out later this morning. Killing observation at PT for the time being as antenna is unpointable. (C134727) 1311 PT tech on site to work on AZ motor 1345 New operator is Alan Kerr 1845-1847 *PT out of point. power glitch. COM 100 2 0000 Date is 2014MAY26/146. 0130 New operator: T. Perreault 0220-0321 *PT, MK out for USNO observation. In an attempt to SCH 200 122 get a good USNO run, will stop this observation before the USNO starts so that the wideband system launches the USNO observation from an idle state -- hopefully this will take care of the T450 switches and DBE errors. 0503-0508 *KP autostowed, power glitch, repointed COM 100 5 0901 %PT wea: light rain 1340 New operator is Alan Kerr 2145 New Operator: Paul Padilla 0000 Date is 2014MAY27/147. 0547 Operator is A. Sowinski. 0645-0750 *PT,MK removed for USNO N4147. SCH 200 130 1329-1333 *LA rotation position error. FRM 100 4 1330-1336 %PT formatter FIFO errors. Won't affect astronomical FMTR 100 6 data but might affect legacy pcal data. 1345 New operator is Alan Kerr 1354-1406 *KP zero-length-error. cleared after 2 scans. REC 100 12 1432-1434 *OV out of point. power glitch. COM 100 2 2055-2100 *OV TimestampSyncError for RDBE-1. TIME 100 5 2100 End. Downtime Summary: Total downtime : 758 min Percentage downtime of observing: 2.8% Average downtime per hour : 1.7 min Total scheduled observing time (# Antennas): 27013 min (7) Notes: * = Entries where data was affected. % = Entries where data may have been affected. & = Entries where the site tech was called out. WEA = Weather entries. MF# = Maintenance form or major downtime category associated with a problem. %AD = The percentage of an antenna affected by a problem. AMD = Total antenna-minutes downtime for a problem. Tsys = System Temperature (TP/SP x Tcal/2) ACU = Antenna Control Unit FRM = Focus/Rotation Mount RFI = Radio Frequency Interference VME = Site control computer CB = Circuit Breaker vclock = Program that compares site clock time to a standard.